• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 11:58:53 on July 31, 2006

    Entry number 176513

    This entry is a followup to: 176510

    keyword=Done with controlled access

    We finished up with what we wanted to do in the Hall. These items are fixed:
    -- LHRS bogie control has been repaired
    -- Radiator camera has a telephoto lens added and was moved to the back
    of the back wall
    -- 'P-Touch' labels applied over Ron's marked radiator positions for
    better visibility

    Items still pending:

    Rick Gonzales took a look at the 1HA1H03B Harp scanner. He found the
    problem and requires a 'custom' part to fix it. It will be a day or two
    to come up with a fix.

    Beam calorimeter chiller repair (Arne F.) is waiting on a part. Not here
    as of this morning, but should be here soon.

    Final fix of the Hall A position FFB issue requires beeam (Dixon).